# T2VEval: Benchmark Dataset and Objective Evaluation Method for T2V-generated Videos

### v5

@@ III-B Compilation and Screening of Prompts
Referring to the scheme adopted by [feng2024tc, liu2024evalcrafter, huang2024vbench], we utilized a large language model (LLM) to pre-generate prompts, followed by manual refinement, as illustrated in Figure 3. Basic elements for each category were predefined and organized into JSON files, which were used as inputs for the LLM to generate over 50 candidates per category, each limited to 50 words. The prompts are divided into four categories: human, animal, landscape, and imagination. We finalized 100 test prompts: 30 for human, 14 for animal, 38 for landscape, and 18 for imagination. These prompts are designed to align with the evaluation dimension, covering aspects such as real world scenes, cultural characteristics, special weather, actions, expressions, spatial relationships and visual styles, as detailed in Figure 4(b).
TABLE I: The rankings of the models being evaluated. The red font represents the highest score, the blue font represents the second highest score, and the green font represents the lowest score. *Sora’s scores only includes 48 prompts’ MOSs, which may have certain limitations.
@@
III-C Video Generation
We select 13 most representative and advanced T2V models, including Sora, Pika 1.0, Runway Gen-3 Alpha, Open-Sora 1.2, PixVerse V2, Show-1, VideoCrafter, Dreamina 1.0, Kling 1.0, Dream Machine 1.0, Pixeling, and TypeMovie. Details of these models are provided in Supplementary Table S1. We generated a total of 1,783 videos including 48 publicly available videos from Sora and 1,735 videos generated independently. The distribution of video quantities across each category is illustrated in Figure 4(a).
III-D Collection of Subjective Scores
Subjective evaluations were conducted in accordance with the criteria of the Recommendation ITU-R BT.500 [bt2019500]. The evaluation process was performed in a controlled laboratory setting to maintain consistent video viewing and evaluation conditions. The 5-point continuous quality scale method was employed, with the scores categorized scores as follows: 0-1 (bad), 1-2 (poor), 2-3 (fair), 3-4 (good), and 4-5 (excellent). Each video was assessed by 7 to 9 professional evaluators, with a total of 73 evaluators recruited for the experiment.
$\textit{MOS}_{ijk}=\frac{1}{N}\sum_{l=1}^{N}S_{ijkl}$ (1)
Supplementary Figure S1 shows the MOS distribution for each dimension, demonstrating that the scores cover nearly the entire possible range and follow a normal distribution. This confirms the comprehensiveness and validity of the subjective scores.
III-E Analysis of T2V Model Performance
@@
where $M$ is the number of prompts. Table I provides the MOS scores of the 13 models across all evaluation dimensions, while Supplementary Figure S3 visualizes the score distribution using a violin plot. Notably, Sora achieves the highest scores in all dimensions, exhibiting a highly concentrated and stable distribution. However, it is important to note that Sora’s ranking is based solely on its 48 demo videos, which introduces certain limitations to the evaluation.
@@
Figure 5: Examples of realness distortion in T2V-generated videos. (a) It is impossible to simulate the natural feeling of fish swimming in water. (b) The limbs of the person are distorted and the action of ”catching an apple” is not expressed.
Finally, we analyzed the performance of these models in each category, and the results are shown in Supplementary Figure S2. Runway Gen-3 consistently outperforms other models across all prompt categories, while Open-Sora 1.2, VideoCrafter, and Show-1 exhibit the weakest performance. This aligns with the conclusions presented in Table I. Furthermore, it is evident that regardless of the prompt category, the scores for the realness dimension are generally lower than those of other dimensions. This highlights that the primary challenge faced by current T2V models lies in accurately understanding and representing the objective laws of the real world, including physical principles and cognitive commonsense knowledge. Specifically, these models struggle to accurately understand and represent the physiological structures (e.g., facial features, limbs), movement patterns (e.g., running, dancing), and physical laws (e.g., sliding, dripping) of organisms in the real world, as illustrated in Figure 5.
IV T2VEval: a Multi-Branch Fusion Scheme for Objective Evaluation
@@
IV-A Technical Quality Encoder
The technical quality encoder aims to represent distortions in both the spatial and temporal domains, including noise, overexposure/underexposure, motion blur, jitter etc. This part can be understood as a conventional video quality assessment (VQA) problem. Recent studies have demonstrated that the Swin-T 3D architecture is well-suited for a wide range of VQA tasks [wu2022fast, wu2023exploring]. Leveraging its 3D sliding window attention mechanism, the model excels at extracting spatio-temporal features, enabling effective analysis of video distortions across both spatial and temporal domains. Therefore, we directly adopt the Tec branch from DOVER [wu2023exploring] as the backbone of our technical quality encoder. Given a video clip $v$, we can extract its corresponding technical quality feature $f_{q}$:
[…]
IV-B Realness Encoder
@@
Inspired by natural language processing work such as BERT [devlin2018bert] and PPT [gu2021ppt], we designed a pre-training task for the realness evaluation, named AI-generated videos detection. We constructed a pre-training dataset comprising 5,000 AI-generated videos and 5,000 human-made videos, sourced from the DeMemba datasets [chen2024demamba]. Using this dataset, we trained a classification network with ConvNeXt 3D [liu2022convnet] as the backbone architecture. This network achieves a classification accuracy of 98% on the training dataset and an 83% zero-shot classification accuracy on the T2VEval-Bench dataset. We believe that the AI-generated video detection task and the T2V-generated video realness evaluation task exhibit stronger correlations. Compared to conventional ImageNet-based pre-training method, our pre-training strategy enables the model to better understand and characterize the realness flaws in T2V-generated videos. For video clip $v$, the realness encoder can extract the corresponding realness feature $f_{r}$:
@@
IV-C Text-Video Alignment Encoder
Text-video alignment refers to the consistency between the video content and the text prompt. CLIP [radford2021learning] and BLIP [li2022blip] are models with strong zero-shot text-to-image matching capabilities. However, experiments [chivileva2023measuring, huang2024vbench, liu2024evalcrafter, liu2024fetv] have shown that simply using CLIP or BLIP has a low correlation with the subjective scores. Additionally, some approaches attempt to generate captions for the video using MLLM (Multimodal Large Language Models) and then compare the similarity between the captions and prompts [chivileva2023measuring], but results indicate this method is also unreliable.
@@ IV-C Text-Video Alignment Encoder
[…]
IV-D Integrated Scoring prediction Module
The proposed Integrated Scoring Regression Module consists of three key components: the vision fidelity fusion unit, the integrated feature fusion unit, and the scoring prediction head. Inspired by T2VQA [wu2024towards], the module adopts a Transformer-based architecture to achieve a progressive and effective integration of features.
@@ IV-D Integrated Scoring prediction Module
LLM has achieved excellent results on IQA and VQA tasks [wu2024q, wu2023q, wu2025towards, zhang2024q]. Inspired by this, we employed LLM as the scoring prediction head in T2VEval and guided LLM by designing a special prompt ”Please rate the quality of the video”. The output of LLM is the probability distribution of the five quality levels (bad, poor, fair, good, and excellent). We assigned weights of 1-5 to these five quality levels. Finally, the predicted score was obtained by calculating the softmax of each token and multiplying it by the corresponding weight.
@@
TABLE II: Performance comparison of different models. Red: the best
@@ IV-E Training Strategy
To further enhance performance, we adopted a dual-loss optimization strategy [li2022blindly] for each branch. This strategy combines Pearson’s Linear Correlation Coefficient (PLCC) and ranking loss, ensuring robust optimization tailored to each evaluation dimension.
@@
V Experiments and results
V-A Implement Details
@@
For the text-video alignment branch, we used the BLIP encoder [li2022blip] to extract features, while the technical quality encoder was initialized with weights pre-trained by Swin-T on the Kinetics-400 dataset [kay2017kinetics]. For the realness encoder, we initialized it with pre-trained parameters on a self-built classification dataset. The LLM used in the scoring prediction head is the 7B version of Vicuna v1.5 [zheng2023judging], which is fine-tuned from Llama2 [touvron2023llama]. What needs illustration is that the gradients of the Vicuna v1.5 were frozen during training.
@@
V-B Performance Comparison
We utilized CLIPSim [wu2021godiva], BLIP [li2022blip], ImageReward [xu2024imagereward], UMTScore [liu2024fetv], Simple VQA [sun2022deep], Fast-VQA [wu2022fast], DOVER [wu2023exploring], and T2VQA [wu2024towards] as reference algorithms for performance comparison experiments. CLIPSim computes the average similarity between text and individual video frames. BLIP and ImageReward are adapted to convert image quality assessment (IQA) metrics into VQA metrics, while UMTScore is tailored to evaluate text-video alignment. T2VQA is specifically designed for T2V evaluation. DOVER, Simple VQA, and Fast-VQA are effective models for VQA. We employed pre-trained weights for CLIPSim, BLIP, ImageReward, and UMTScore for zero-shot testing, while Simple VQA, Fast-VQA, DOVER, and T2VQA are fine-tuned on the T2VEval-Bench.
Table II compares the performance of T2VEval with other methods. The results demonstrate that our method achieves state-of-the-art performance on T2VEval-Bench, surpassing the second-best method by nearly 4%. The performance of zero-shot models is notably poor, which may be attributed to their sole focus on text-video alignment and inability to capture temporal information between video frames. Furthermore, these methods are typically trained and tested on datasets with human-made videos, overlooking the domain gap between AI-generated videos and human-made videos. Although VQA models achieve higher scores, they still fall significantly short of T2VQA and T2VEval, further underscoring the necessity of multi-branch fusion evaluation framework evaluation frameworks.
@@ V-B Performance Comparison
Additionally, we fine-tuned and tested T2VEval on the T2VQA-DB dataset, following the same data partitioning method as T2VEval-Bench. The results, averaged across 10 partitions, are also presented in Table II. These experiments confirm that T2VEval also achieves state-of-the-art performance in this dataset.
@@
V-C Ablation Studies
To systematically evaluate the contribution of core architectural components in T2VEval, we conducted comprehensive ablation studies. Specifically, in Table III, we conducted experiments to evaluate (i) different vision fidelity fusion mechanisms and (ii) differential pre-training strategies applied to the realness encoder. Additionally, in Table IV, we explored the contribution and effect of each branch, aiming to gain a deeper understanding of how these components interact and influence the overall performance of T2VEval.
V-C1 Ablation Experiment of Vision Fidelity Fusion Strategy
T2VEval employs a cross-attention mechanism to fuse technical quality and realness features. In previous studies, alternative fusion methods, such as feature summation and concatenation, have also been utilized. To evaluate their effectiveness, we compared the cross-attention fusion strategy against addition and concatenation methods (Addition is denoted as T2VEval-Add, and concatenation is denoted as T2VEval-Cat). As shown in Table III, the cross-attention method significantly outperforms the other two approaches.
This superiority can be explained by the dynamic nature of cross-attention, which assigns adaptive weights to different features based on their intrinsic relevance. Unlike summation and concatenation, cross-attention selectively highlights valuable information while maintaining contextual integrity. Feature summation, on the other hand, may result in information loss due to the ”smoothing” effect during addition, which reduces the distinctions between features. While concatenation preserves all feature information, it fails to establish meaningful interactions between features.
@@ V-C1 Ablation Experiment of Vision Fidelity Fusion Strategy
Results are shown in Table III. These results may be due to the fact that realness and technical quality are completely dependent on visual content and have a stronger task-dependence. In contrast, due to the existence of the text modality, fusing visual-related features (whether realness feature or technical quality feature) with consistency features first may lead to interference among various modal features.
TABLE III: Performance comparison of ablation studies. Red: the best
@@
TABLE IV: Ablation study on the usage of technical quality branch, realness branch and alignment branch in T2VEval. Red: the best
@@
V-C2 Ablation Experiment of Realness Branch Pre-train Strategy
When designing the realness evaluation branch, we introduced the pre-training task of AI-generated video detection, instead of simply using ImageNet pre-trained weights or random initialization. We replaced the pre-trained weights of Conv3D with ImageNet weights (denoted as T2VEval-ImageNet) and random initialization (denoted as T2VEval-Random) respectively, and fine-tuned them in the same training environment. The experimental results are shown in Table III. In addition, we further used Grad-CAM to analyze the impact of different pre-training strategies on the model’s focus area, as shown in Figure 8. The figure demonstrates the activation maps of the 3D network loaded with different pre-trained weights.

### v6

@@ 3.2 Compilation and Screening of Prompts
Referring to the scheme adopted by Feng et al. (2024); Liu et al. (2024a); Huang et al. (2024), we utilized a large language model (LLM) to pre-generate prompts, followed by manual refinement, as illustrated in Figure 3. Basic elements for each category were predefined and organized into JSON files, which were used as inputs for the LLM to generate over 50 candidates per category, each limited to 50 words. The prompts are divided into four categories: human, animal, landscape, and imagination. We finalized 100 test prompts: 30 for human, 14 for animal, 38 for landscape, and 18 for imagination. These prompts are designed to align with the evaluation dimension, covering aspects such as real world scenes, cultural characteristics, special weather, actions, expressions, spatial relationships and visual styles, as detailed in Figure 4(b).
3.3 Video Generation
We select 13 most representative and advanced T2V models, including Sora, Pika 1.0, Runway Gen-3 Alpha, Open-Sora 1.2, PixVerse V2, Show-1, VideoCrafter, Dreamina 1.0, Kling 1.0, Dream Machine 1.0, Pixeling, and TypeMovie. Details of these models are provided in Table 1. We generated a total of 1,783 videos including 48 publicly available videos from Sora and 1,735 videos generated independently. The distribution of video quantities across each category is illustrated in Figure 4(a).
Figure 5: Distribution of subjective scores in each dimension. (a) The overall impression MOS distribution. (b) The realness MOS distribution. (c) The overall text-video consistency MOS distribution. (d) The technical quality MOS distribution.
Figure 6: The Violin distribution maps of the model’s subjective score in each evaluation dimension
Figure 7: The radar maps of each model under different prompt’s categories. Considering that Sora only had 48 demo videos, which may affect the reliability of evaluation, we did not include it in our analysis.
3.4 Collection of Subjective Scores
Subjective evaluations were conducted in accordance with the criteria of the Recommendation ITU-R BT.500 BT.500 (2023). The evaluation process was performed in a controlled laboratory setting to maintain consistent video viewing and evaluation conditions. The 5-point continuous quality scale method was employed, with the scores categorized scores as follows: 0-1 (bad), 1-2 (poor), 2-3 (fair), 3-4 (good), and 4-5 (excellent). Each video was assessed by 7 to 9 professional evaluators, with a total of 73 evaluators recruited for the experiment.
$\textit{MOS}_{ijk}=\frac{1}{N}\sum_{l=1}^{N}S_{ijkl}$ (1)
Figure 5 shows the MOS distribution for each dimension, demonstrating that the scores cover nearly the entire possible range and follow a normal distribution. This confirms the comprehensiveness and validity of the subjective scores.
Table 2: The rankings of the models being evaluated. The red font represents the highest score, the blue font represents the second highest score, and the green font represents the lowest score. *Sora’s scores only includes 48 prompts’ MOSs, which may have certain limitations.
@@ 3.4 Collection of Subjective Scores
3.5 Analysis of T2V Model Performance
@@
where $M$ is the number of prompts. Table 2 provides the MOS scores of the 13 models across all evaluation dimensions, while Figure 6 visualizes the score distribution using a violin plot. Notably, Sora achieves the highest scores in all dimensions, exhibiting a highly concentrated and stable distribution. However, it is important to note that Sora’s ranking is based solely on its 48 demo videos, which introduces certain limitations to the evaluation.
@@
Figure 8: Examples of realness distortion in T2V-generated videos. (a) It is impossible to simulate the natural feeling of fish swimming in water. (b) The limbs of the person are distorted and the action of "catching an apple" is not expressed.
Finally, we analyzed the performance of these models in each category, and the results are shown in Figure 7. Runway Gen-3 consistently outperforms other models across all prompt categories, while Open-Sora 1.2, VideoCrafter, and Show-1 exhibit the weakest performance. This aligns with the conclusions presented in Table 2. Furthermore, it is evident that regardless of the prompt category, the scores for the realness dimension are generally lower than those of other dimensions. This highlights that the primary challenge faced by current T2V models lies in accurately understanding and representing the objective laws of the real world, including physical principles and cognitive commonsense knowledge. Specifically, these models struggle to accurately understand and represent the physiological structures (e.g., facial features, limbs), movement patterns (e.g., running, dancing), and physical laws (e.g., sliding, dripping) of organisms in the real world, as illustrated in Figure 8.
4 T2VEval: a Multi-Branch Fusion Scheme for Objective Evaluation
@@
4.1 Technical Quality Encoder
The technical quality encoder aims to represent distortions in both the spatial and temporal domains, including noise, overexposure/underexposure, motion blur, jitter etc. This part can be understood as a conventional video quality assessment (VQA) problem. Recent studies have demonstrated that the Swin-T 3D architecture is well-suited for a wide range of VQA tasks Wu et al. (2022b, 2023a). Leveraging its 3D sliding window attention mechanism, the model excels at extracting spatio-temporal features, enabling effective analysis of video distortions across both spatial and temporal domains. Therefore, we directly adopt the Tec branch from DOVER Wu et al. (2023a) as the backbone of our technical quality encoder. Given a video clip $v$, we can extract its corresponding technical quality feature $f_{q}$:
$f_{q}~=~Swin3D(v)$ (3)
4.2 Realness Encoder
@@
Inspired by natural language processing work such as BERT Devlin (2018) and PPT Gu et al. (2021), we designed a pre-training task for the realness evaluation, named AI-generated videos detection. We constructed a pre-training dataset comprising 5,000 AI-generated videos and 5,000 human-made videos, sourced from the DeMemba datasets Chen et al. (2024a). Using this dataset, we trained a classification network with ConvNeXt 3D Liu et al. (2022a) as the backbone architecture. This network achieves a classification accuracy of 98% on the training dataset and an 83% zero-shot classification accuracy on the T2VEval-Bench dataset. We believe that the AI-generated video detection task and the T2V-generated video realness evaluation task exhibit stronger correlations. Compared to conventional ImageNet-based pre-training method, our pre-training strategy enables the model to better understand and characterize the realness flaws in T2V-generated videos. For video clip $v$, the realness encoder can extract the corresponding realness feature $f_{r}$:
@@
4.3 Text-Video Alignment Encoder
Text-video alignment refers to the consistency between the video content and the text prompt. CLIP Radford et al. (2021) and BLIP Li et al. (2022b) are models with strong zero-shot text-to-image matching capabilities. However, experiments Chivileva et al. (2023); Huang et al. (2024); Liu et al. (2024a, b) have shown that simply using CLIP or BLIP has a low correlation with the subjective scores. Additionally, some approaches attempt to generate captions for the video using MLLM (Multimodal Large Language Models) and then compare the similarity between the captions and prompts Chivileva et al. (2023), but results indicate this method is also unreliable.
@@ 4.3 Text-Video Alignment Encoder
$f_{c}=~SA(CA\{BLIP(v_{i}),BLIP(t)\}_{i=1}^{N})$ (5)
4.4 Integrated Scoring prediction Module
The proposed Integrated Scoring Regression Module consists of three key components: the vision fidelity fusion unit, the integrated feature fusion unit, and the scoring prediction head. Inspired by T2VQA Wu et al. (2024b), the module adopts a Transformer-based architecture to achieve a progressive and effective integration of features.
@@ 4.4 Integrated Scoring prediction Module
LLM has achieved excellent results on IQA and VQA tasks Wu et al. (2024a, 2023b, 2025); Zhang et al. (2024d). Inspired by this, we employed LLM as the scoring prediction head in T2VEval and guided LLM by designing a special prompt "Please rate the quality of the video". The output of LLM is the probability distribution of the five quality levels (bad, poor, fair, good, and excellent). We assigned weights of 1-5 to these five quality levels. Finally, the predicted score was obtained by calculating the softmax of each token and multiplying it by the corresponding weight.
@@
Table 3: Performance comparison of different models. Red: the best
@@ 4.5 Training Strategy
To further enhance performance, we adopted a dual-loss optimization strategy Li et al. (2022a) for each branch. This strategy combines Pearson’s Linear Correlation Coefficient (PLCC) and ranking loss, ensuring robust optimization tailored to each evaluation dimension.
@@
5 Experiments and results
5.1 Implement Details
@@
For the text-video alignment branch, we used the BLIP encoder Li et al. (2022b) to extract features, while the technical quality encoder was initialized with weights pre-trained by Swin-T on the Kinetics-400 dataset Kay et al. (2017). For the realness encoder, we initialized it with pre-trained parameters on a self-built classification dataset. The LLM used in the scoring prediction head is the 7B version of Vicuna v1.5 Zheng et al. (2023), which is fine-tuned from Llama2 Touvron et al. (2023). What needs illustration is that the gradients of the Vicuna v1.5 were frozen during training.
@@
5.2 Performance Comparison
We utilized CLIPSim Wu et al. (2021), BLIP Li et al. (2022b), ImageReward Xu et al. (2024), UMTScore Liu et al. (2024b), Simple VQA Sun et al. (2022), Fast-VQA Wu et al. (2022b), DOVER Wu et al. (2023a), and T2VQA Wu et al. (2024b) as reference algorithms for performance comparison experiments. CLIPSim computes the average similarity between text and individual video frames. BLIP and ImageReward are adapted to convert image quality assessment (IQA) metrics into VQA metrics, while UMTScore is tailored to evaluate text-video alignment. T2VQA is specifically designed for T2V evaluation. DOVER, Simple VQA, and Fast-VQA are effective models for VQA. We employed pre-trained weights for CLIPSim, BLIP, ImageReward, and UMTScore for zero-shot testing, while Simple VQA, Fast-VQA, DOVER, and T2VQA are fine-tuned on the T2VEval-Bench.
Table 3 compares the performance of T2VEval with other methods. The results demonstrate that our method achieves state-of-the-art performance on T2VEval-Bench, surpassing the second-best method by nearly 4%. The performance of zero-shot models is notably poor, which may be attributed to their sole focus on text-video alignment and inability to capture temporal information between video frames. Furthermore, these methods are typically trained and tested on datasets with human-made videos, overlooking the domain gap between AI-generated videos and human-made videos. Although VQA models achieve higher scores, they still fall significantly short of T2VQA and T2VEval, further underscoring the necessity of multi-branch fusion evaluation framework evaluation frameworks.
@@ 5.2 Performance Comparison
Additionally, we fine-tuned and tested T2VEval on the T2VQA-DB dataset, following the same data partitioning method as T2VEval-Bench. The results, averaged across 10 partitions, are also presented in Table 3. These experiments confirm that T2VEval also achieves state-of-the-art performance in this dataset.
5.3 Ablation Studies
To systematically evaluate the contribution of core architectural components in T2VEval, we conducted comprehensive ablation studies. Specifically, in Table 4, we conducted experiments to evaluate (i) different vision fidelity fusion mechanisms and (ii) differential pre-training strategies applied to the realness encoder. Additionally, in Table 5, we explored the contribution and effect of each branch, aiming to gain a deeper understanding of how these components interact and influence the overall performance of T2VEval.
5.3.1 Ablation Experiment of Vision Fidelity Fusion Strategy
T2VEval employs a cross-attention mechanism to fuse technical quality and realness features. In previous studies, alternative fusion methods, such as feature summation and concatenation, have also been utilized. To evaluate their effectiveness, we compared the cross-attention fusion strategy against addition and concatenation methods (Addition is denoted as T2VEval-Add, and concatenation is denoted as T2VEval-Cat). As shown in Table 4, the cross-attention method significantly outperforms the other two approaches.
@@ 5.3.1 Ablation Experiment of Vision Fidelity Fusion Strategy
This superiority can be explained by the dynamic nature of cross-attention, which assigns adaptive weights to different features based on their intrinsic relevance. Unlike summation and concatenation, cross-attention selectively highlights valuable information while maintaining contextual integrity. Feature summation, on the other hand, may result in information loss due to the "smoothing" effect during addition, which reduces the distinctions between features. While concatenation preserves all feature information, it fails to establish meaningful interactions between features.
@@
Results are shown in Table 4. These results may be due to the fact that realness and technical quality are completely dependent on visual content and have a stronger task-dependence. In contrast, due to the existence of the text modality, fusing visual-related features (whether realness feature or technical quality feature) with consistency features first may lead to interference among various modal features.
Table 4: Performance comparison of ablation studies. Red: the best
@@
Table 5: Ablation study on the usage of technical quality branch, realness branch and alignment branch in T2VEval. Red: the best
@@
5.3.2 Ablation Experiment of Realness Branch Pre-train Strategy
When designing the realness evaluation branch, we introduced the pre-training task of AI-generated video detection, instead of simply using ImageNet pre-trained weights or random initialization. We replaced the pre-trained weights of Conv3D with ImageNet weights (denoted as T2VEval-ImageNet) and random initialization (denoted as T2VEval-Random) respectively, and fine-tuned them in the same training environment. The experimental results are shown in Table 4. In addition, we further used Grad-CAM to analyze the impact of different pre-training strategies on the model’s focus area, as shown in Figure 11. The figure demonstrates the activation maps of the 3D network loaded with different pre-trained weights.